# Spatial Policy: Guiding Visuomotor Robotic Manipulation with Spatial-Aware Modeling and Reasoning

### v1

@@ Appendix A Inference Pipeline
$\mathbf{z}_{\text{coord}}\leftarrow\texttt{EncodeCoord}(\mathbf{p}_{\text{ee}},\mathbf{p}_{\text{obj}}),\quad\mathcal{A}_{1:H}\leftarrow\pi(\mathbf{V}_{1:T},\mathbf{z}_{\text{coord}})$ ;
@@
    $\mathbf{z}_{\text{coord}}\leftarrow\texttt{EncodeCoord}(\mathbf{p}_{\text{ee}},\mathbf{p}_{\text{obj}}),\quad\mathcal{A}_{1:H}\leftarrow\pi(\mathbf{V}_{1:T},\mathbf{z}_{\text{coord}})$ ;
@@
Appendix B Training Details: Video Generation and Action Prediction
Video Generation
@@ Appendix B Training Details: Video Generation and Action Prediction
All models are trained on $128\times 128$ RGB frames sampled as 8-frame clips. We use the AdamW optimizer (Loshchilov and Hutter 2019) with an initial learning rate of $1\times 10^{-4}$ and a cosine learning rate schedule. Training is conducted for 100,000 steps, unless otherwise specified. The standard batch size is 8 for training and 16 for validation, with gradient accumulation set to 1. Mixed precision training is employed using FP16 and automatic mixed precision (AMP). A linear diffusion noise schedule is used with $\beta_{1}=1\times 10^{-4}$ and $\beta_{T}=0.02$ over 1000 steps. Exponential Moving Average (EMA) with a decay factor of 0.999 is updated every 10 training steps. Conditional dropout with a probability of 0.1 is applied to all models except the Baseline. Validation uses 1 or 2 samples per batch depending on the script. Unless otherwise specified, all training datasets are collected and generated from our own single successful executions. All experiments are conducted on a single NVIDIA RTX 4090 GPU (80GB memory), with full-precision (FP32) training completed within 2 days.
@@
Unlike the original Diffusion Policy (Chi et al. 2023), which conditions on past observations and actions, our model relies solely on the current observation, future observation, and spatial context. Specifically, instead of using a standard 3-channel RGB input, we concatenate the current frame and a sampled goal frame to form a 6-channel input image of resolution $128\times 128$, along with spatial coordinates. Given horizon $H=4$, the model predicts $H$ future actions from a single observation step ($T=1$), conditioned on the current frame, spatial coordinates, and a goal frame sampled within the next $K=20$ frames. We adopt a ResNet-50 visual encoder (He et al. 2016) pretrained on ImageNet (Deng et al. 2009), which remains trainable during training. The spatial coordinates are encoded via a MLP, which is also trained end-to-end. The policy network is implemented as a conditional U-Net with diffusion denoising, using a cosine noise schedule ($\beta_{1}=1\times 10^{-4}$, $\beta_{T}=2\times 10^{-2}$) over 100 steps. The model is optimized using AdamW with learning rate $\eta=1\times 10^{-4}$, weight decay $1\times 10^{-6}$, and batch size $B=512$ for both training and validation. The learning rate follows a cosine annealing schedule with $500$ warm-up steps. Exponential Moving Average (EMA) is applied with a maximum decay rate of $0.9999$ to stabilize training. All experiments are conducted on a single NVIDIA RTX 4090 GPU (40GB memory), with full-precision (FP32) training completed within 3 days.
@@
Appendix C Tasks in Meta-World
In the Meta-World benchmark (Yu et al. 2020), we evaluated the performance on 11 representative robotic manipulation tasks shown in Figure 5, which are designed to assess various manipulation skills and generalization abilities of reinforcement learning algorithms. These tasks are categorized by their difficulty levels as follows:
@@ Appendix C Tasks in Meta-World
Figure 5: Examples of 11 tasks in meta-world
Appendix D Dataset Construction
We construct our dataset from expert demonstrations collected in the Meta-World (Yu et al. 2020) simulator. For each task, successful execution trajectories are recorded and stored as NPZ and PKL files. Each trajectory contains RGB, depth, and mask images at each timestep, along with task metadata such as environment name, frame index, and camera parameters.
@@ Appendix E VLM Prompt Design
    Spatial plan table generation: As described in Section  D, we format spatial plan table into structured multi-step plans of the form:
    1Plan:
@@ Conclusion:
Appendix F Frame Matching Mechanism
@@
where $\text{Sim}_{\text{geo}}$ is the similarity based on contour and shape features (Lowe 2004), $\text{Sim}_{\text{pos}}$ encodes block-wise intensity statistics and center-of-mass shifts, SSIM is the structural similarity index between grayscale frames (Wang et al. 2004), and $\text{Sim}_{\text{flow}}$ measures the inverse magnitude of optical flow (Farnebäck 2003). The weights are set to $w_{\text{geo}}{=}0.35$, $w_{\text{pos}}{=}0.35$, $w_{\text{ssim}}{=}0.2$, and $w_{\text{flow}}{=}0.1$. A frame is considered matched if:
@@
Appendix G Robustness to Occlusions via Masking
To evaluate the robustness of our video generation model under partial visual occlusions, we introduce a masking mechanism before feeding the current frame into the diffusion model. Specifically, we randomly mask a portion of the image input, simulating occlusions such as those caused by environmental factors or sensor noise.

### v2

@@ A.1 Inference Pipeline
$\mathbf{z}_{\text{coord}}\leftarrow\texttt{EncodeCoord}(\mathbf{p}_{\text{ee}},\mathbf{p}_{\text{obj}}),\qquad\mathcal{A}_{1:H}\leftarrow\pi(\mathbf{V}_{1:T},\mathbf{z}_{\text{coord}})$ ;
@@
    $\mathbf{z}_{\text{coord}}\leftarrow\texttt{EncodeCoord}(\mathbf{p}_{\text{ee}},\mathbf{p}_{\text{obj}}),\qquad\mathcal{A}_{1:H}\leftarrow\pi(\mathbf{V}_{1:T},\mathbf{z}_{\text{coord}})$ ;
@@
A.2 Training Details: Video Generation and Action Prediction
A.2.1 Video Generation
@@
All models are trained on $128\times 128$ RGB frames sampled as 8-frame clips. We use the AdamW optimizer [adam] with an initial learning rate of $1\times 10^{-4}$ and a cosine learning rate schedule. Training is conducted for 100,000 steps, unless otherwise specified. The standard batch size is 8 for training and 16 for validation, with gradient accumulation set to 1. Mixed precision training is employed using FP16 and automatic mixed precision (AMP). A linear diffusion noise schedule is used with $\beta_{1}=1\times 10^{-4}$ and $\beta_{T}=0.02$ over 1000 steps. Exponential Moving Average (EMA) with a decay factor of 0.999 is updated every 10 training steps. Conditional dropout with a probability of 0.1 is applied to all models except the Baseline. Validation uses 1 or 2 samples per batch depending on the script. Unless otherwise specified, all training datasets are collected and generated from our own single successful executions. All experiments are conducted on a single NVIDIA RTX 4090 GPU (80GB memory), with full-precision (FP32) training completed within 2 days.
@@ A.2.2 Action Prediction
Unlike the original Diffusion Policy [diffusionpolicy], which conditions on past observations and actions, our model relies solely on the current observation, future observation, and spatial context. Specifically, instead of using a standard 3-channel RGB input, we concatenate the current frame and a sampled goal frame to form a 6-channel input image of resolution $128\times 128$, along with spatial coordinates. Given horizon $H=4$, the model predicts $H$ future actions from a single observation step ($T=1$), conditioned on the current frame, spatial coordinates, and a goal frame sampled within the next $K=20$ frames. We adopt a ResNet-50 visual encoder [resnet] pretrained on ImageNet [imagenet], which remains trainable during training. The spatial coordinates are encoded via a MLP, which is also trained end-to-end. The policy network is implemented as a conditional U-Net with diffusion denoising, using a cosine noise schedule ($\beta_{1}=1\times 10^{-4}$, $\beta_{T}=2\times 10^{-2}$) over 100 steps. The model is optimized using AdamW with learning rate $\eta=1\times 10^{-4}$, weight decay $1\times 10^{-6}$, and batch size $B=512$ for both training and validation. The learning rate follows a cosine annealing schedule with $500$ warm-up steps. Exponential Moving Average (EMA) is applied with a maximum decay rate of $0.9999$ to stabilize training. All experiments are conducted on a single NVIDIA RTX 4090 GPU (40GB memory), with full-precision (FP32) training completed within 3 days.
@@
A.3 Tasks in Meta-World and iThor
Meta-World. In the Meta-World benchmark [metaworld], we evaluated the performance on 11 representative robotic manipulation tasks shown in Figure 8, which are designed to assess various manipulation skills and generalization abilities of reinforcement learning algorithms. These tasks are categorized by their difficulty levels as follows:
@@ A.3 Tasks in Meta-World and iThor
iThor. To assess generalization in visually diverse, embodied environments, we further evaluate on 12 navigation-style manipulation tasks in iTHOR, spanning four rooms: Kitchen, Living Room, Bedroom, and Bathroom. Tasks include locating or approaching objects under varying viewpoints and scene layouts, such as Blinds, Bread, DeskLamp, Laptop, Mirror, Painting, Pillow, SoapBar, Spatula, Television, Toaster, and ToiletPaper. The full set of 4 rooms and 12 tasks is shown in Figure 9.
Together, these two suites provide broad coverage of manipulation and navigation challenges, enabling a comprehensive evaluation of visuomotor generalization across simulated robotic arms and interactive 3D household environments.
Figure 8: Generated video plans on the $1$ Meta-World tasks.
Figure 9: Generated video plans across 12 tasks in four iTHOR room types (Kitchen, Living Room, Bedroom, Bathroom).
A.4 Dataset Construction
We construct our dataset from expert demonstrations collected in the Meta-World [metaworld] simulator. For each task, successful execution trajectories are recorded and stored as NPZ and PKL files. Each trajectory contains RGB, depth, and mask images at each timestep, along with task metadata such as environment name, frame index, and camera parameters.
@@ A.5 VLM Prompt Design
    Spatial plan table generation: As described in Section  A.4, we format spatial plan table into structured multi-step plans of the form:
    Plan:
@@ Conclusion:
A.6 Frame Matching Mechanism
@@
where $\text{Sim}_{\text{geo}}$ is the similarity based on contour and shape features [lowe2004distinctive], $\text{Sim}_{\text{pos}}$ encodes block-wise intensity statistics and center-of-mass shifts, SSIM is the structural similarity index between grayscale frames [wang2004image], and $\text{Sim}_{\text{flow}}$ measures the inverse magnitude of optical flow [farneback2003two]. The weights are set to $w_{\text{geo}}{=}0.35$, $w_{\text{pos}}{=}0.35$, $w_{\text{ssim}}{=}0.2$, and $w_{\text{flow}}{=}0.1$. A frame is considered matched if: